# High Probability Bound for Cross-Learning Contextual Bandits with Unknown Context Distributions

### v1

@@ B.3 Upper Bounding $\operatorname{\mathbf{bias}}_{4}$
We then bound the forth term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\right\rangle$. Similar to the previous analysis, we decompose it as follows:
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
@@
$\sum_{e}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle.$
@@
$=$ $\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
$+\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c}-\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right].$
@@
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c}-\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
is bounded by $\frac{98KT\iota}{L}+\frac{\gamma^{2}LKT}{\iota}.$ Furthermore, condition on $\operatorname{\mathcal{H}}_{e-1}$, the indicator function $F_{e}$ is effected only by randomness within time steps $t\in\operatorname{\mathcal{T}_{e}^{f}}$, thus the indicator function $F_{e}$ is conditional independent with the probability vector $\tilde{p}_{t,c}$. We have
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
$=$ $\sum_{e}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\mathbb{E}\left[\tilde{p}_{t,c}\big{|}\operatorname{\mathcal{H}}_{e-1}\right],\mathbb{E}\left[(\tilde{\ell}_{t,c}-\hat{\ell}_{t,c})F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]\right\rangle$
$\leq$ $\sum_{e}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}\frac{\tilde{p}_{t,c}(a)\gamma}{f_{e}(a)}.$ (Lemma 5)
By Lemma 4, whenever event $G$ holds, the ratio $\frac{\tilde{p}_{t,c}(a)}{f_{e}(a)}\leq 4.$ Thus we have
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]\leq 4\gamma KT$
@@
We further have $\left|\left\langle\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\right|\leq\frac{1}{\gamma}$. Thus we have
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}(L_{e}-1)\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
@@
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\tilde{p}_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
@@
$=$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}$
$=$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}p_{t,c}(a)f_{e}(a)\ell_{t,c}(a)\left(\frac{1}{f_{e}(a)+\gamma}-\frac{1}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}\right)F_{e}L_{e}$
$=$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}\tilde{p}_{t,c}(a)f_{e}(a)\ell_{t,c}(a)\frac{\hat{f}_{e}(a)-f_{e}(a)+\frac{1}{2}\gamma}{(f_{e}(a)+\gamma)(\hat{f}_{e}(a)+\frac{3}{2}\gamma)}F_{e}L_{e}.$
@@
$\leq$ $\left|\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}\tilde{p}_{t,c}(a)f_{e}(a)\ell_{t,c}(a)\frac{\hat{f}_{e}(a)-f_{e}(a)+\frac{1}{2}\gamma}{(f_{e}(a)+\gamma)(\hat{f}_{e}(a)+\frac{3}{2}\gamma)}\right|F_{e}L_{e}$
$\leq$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{a}\tilde{p}_{t}(a)\left|\frac{\hat{f}_{e}(a)-f_{e}(a)+\frac{1}{2}\gamma}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}\right|F_{e}L_{e}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle.$
@@
$J_{t}\triangleq\operatorname{\mathbbm{1}}\left(\forall a,\tilde{p}_{t}(a)\leq 4f_{e}(a)\right).$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle(1-F_{e}J_{t}).$
Since the auxiliary probability vector $\tilde{p}_{t,c}$ is determined at time $t-1$, the indicator function $J_{t}$ is also determined at time $t-1$. Furthermore, the indicator function $F_{e}$ is determined at epoch $e-1$. Thus the product of indicator functions $F_{e}J_{t}$ is measurable under filtration $\operatorname{\mathcal{H}}_{t-1}$. We have
$\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
$=$ $F_{e}J_{t}\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},(\tilde{\ell}_{t,c}-\hat{\ell}_{t,c})-\mathbb{E}\left[(\tilde{\ell}_{t,c}-\hat{\ell}_{t,c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
@@
$\left\{\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$
@@
$\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}$
@@
$\left|\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\right|F_{e}J_{t}$
@@
$+\left|\sum_{c}\Pr(c)\sum_{a}p_{t,c}(a)\ell_{t,c}(a)\left(\frac{\operatorname{\mathbbm{1}}(a_{t}=a)}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}-\frac{f_{e}(a)}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}\right)\right|F_{e}J_{t}$
@@
$+\sum_{c}\Pr(c)p_{t,c}(a_{t})\frac{\ell_{t,c}(a_{t})}{\hat{f}_{e}(a_{t})+\frac{3}{2}\gamma}F_{e}J_{t}$
$+\sum_{c}\Pr(c)\sum_{a}p_{t,c}(a)\ell_{t,c}(a)\frac{f_{e}(a)}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}F_{e}J_{t}$
$\leq$ $\left(\frac{p_{t}(a_{t})}{f_{e}(a_{t})+\gamma}+1+\frac{p_{t}(a_{t})}{\hat{f}_{e}(a_{t})+\frac{3}{2}\gamma}+\sum_{a}p_{t}(a)\frac{f_{e}(a)}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}\right)F_{e}J_{t}$
@@
$\left\{\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}},$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}-\mathbb{E}\left[\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\leq 15\sqrt{T\log(\frac{1}{\delta})}$

### v2

@@ B.3 Upper Bounding $\operatorname{\mathbf{bias}}_{4}$
We then bound the forth term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\right\rangle$. Similar to the previous analysis, we decompose it as follows:
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
@@
$\sum_{e}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle.$
@@
$=$ $\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
$+\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c}-\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right].$
@@
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c}-\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
is bounded by $\frac{98KT\iota}{L}+\frac{\gamma^{2}LKT}{\iota}.$ Furthermore, condition on $\operatorname{\mathcal{H}}_{e-1}$, the indicator function $F_{e}$ is effected only by randomness within time steps $t\in\operatorname{\mathcal{T}_{e}^{f}}$, thus the indicator function $F_{e}$ is conditional independent with the probability vector $\widetilde{p}_{t,c}$. We have
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
$=$ $\sum_{e}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\mathbb{E}\left[\widetilde{p}_{t,c}\big{|}\operatorname{\mathcal{H}}_{e-1}\right],\mathbb{E}\left[(\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c})F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]\right\rangle$
$\leq$ $\sum_{e}\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}\frac{\widetilde{p}_{t,c}(a)\gamma}{f_{e}(a)}.$ (Lemma 5)
By Lemma 4, whenever event $G$ holds, the ratio $\frac{\widetilde{p}_{t,c}(a)}{f_{e}(a)}\leq 4.$ Thus we have
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]\leq 4\gamma KT$
@@
We further have $\left|\left\langle\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\right|\leq\frac{1}{\gamma}$. Thus we have
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}(L_{e}-1)\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
@@
$\sum_{e}\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle\widetilde{p}_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}\big{|}\operatorname{\mathcal{H}}_{e-1}\right]$
@@
$=$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}L_{e}$
$=$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}p_{t,c}(a)f_{e}(a)\ell_{t,c}(a)\left(\frac{1}{f_{e}(a)+\gamma}-\frac{1}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\right)F_{e}L_{e}$
$=$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}\widetilde{p}_{t,c}(a)f_{e}(a)\ell_{t,c}(a)\frac{\widehat{f}_{e}(a)-f_{e}(a)+\frac{1}{2}\gamma}{(f_{e}(a)+\gamma)(\widehat{f}_{e}(a)+\frac{3}{2}\gamma)}F_{e}L_{e}.$
@@
$\leq$ $\left|\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{c}\Pr(c)\sum_{a}\widetilde{p}_{t,c}(a)f_{e}(a)\ell_{t,c}(a)\frac{\widehat{f}_{e}(a)-f_{e}(a)+\frac{1}{2}\gamma}{(f_{e}(a)+\gamma)(\widehat{f}_{e}(a)+\frac{3}{2}\gamma)}\right|F_{e}L_{e}$
$\leq$ $\sum_{t\in\operatorname{\mathcal{T}_{e}^{\ell}}}\sum_{a}\widetilde{p}_{t}(a)\left|\frac{\widehat{f}_{e}(a)-f_{e}(a)+\frac{1}{2}\gamma}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\right|F_{e}L_{e}$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle.$
@@
$J_{t}\triangleq\operatorname{\mathbbm{1}}\left(\forall a,\widetilde{p}_{t}(a)\leq 4f_{e}(a)\right).$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle(1-F_{e}J_{t}).$
Since the auxiliary probability vector $\widetilde{p}_{t,c}$ is determined at time $t-1$, the indicator function $J_{t}$ is also determined at time $t-1$. Furthermore, the indicator function $F_{e}$ is determined at epoch $e-1$. Thus the product of indicator functions $F_{e}J_{t}$ is measurable under filtration $\operatorname{\mathcal{H}}_{t-1}$. We have
$\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
$=$ $F_{e}J_{t}\mathbb{E}\left[\sum_{c}\Pr(c)\left\langle p_{t,c},(\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c})-\mathbb{E}\left[(\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\big{|}\operatorname{\mathcal{H}}_{t-1}\right]$
@@
$\left\{\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}}$
@@
$\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}$
@@
$\left|\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle\right|F_{e}J_{t}$
@@
$+\left|\sum_{c}\Pr(c)\sum_{a}p_{t,c}(a)\ell_{t,c}(a)\left(\frac{\operatorname{\mathbbm{1}}(a_{t}=a)}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}-\frac{f_{e}(a)}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\right)\right|F_{e}J_{t}$
@@
$+\sum_{c}\Pr(c)p_{t,c}(a_{t})\frac{\ell_{t,c}(a_{t})}{\widehat{f}_{e}(a_{t})+\frac{3}{2}\gamma}F_{e}J_{t}$
$+\sum_{c}\Pr(c)\sum_{a}p_{t,c}(a)\ell_{t,c}(a)\frac{f_{e}(a)}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}F_{e}J_{t}$
$\leq$ $\left(\frac{p_{t}(a_{t})}{f_{e}(a_{t})+\gamma}+1+\frac{p_{t}(a_{t})}{\widehat{f}_{e}(a_{t})+\frac{3}{2}\gamma}+\sum_{a}p_{t}(a)\frac{f_{e}(a)}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\right)F_{e}J_{t}$
@@
$\left\{\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\right\}_{t\in\operatorname{\mathcal{T}^{\ell}}},$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}-\mathbb{E}\left[\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right\rangle F_{e}J_{t}\leq 15\sqrt{T\log(\frac{1}{\delta})}$